OpenStudy (anonymous):

solve for n 9/8 (fraction)n=-54 Can any of you guys explain this to me please and thankyou

OpenStudy (anonymous):

please

OpenStudy (anonymous):

i need a answer quick!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

@uChezzy

OpenStudy (anonymous):

@uChezzy

OpenStudy (uchezzy):

I_always_Smiling

OpenStudy (anonymous):

can u help me plz

OpenStudy (uchezzy):

@I_always_Smiling

OpenStudy (anonymous):

@I_Always_Smiling

OpenStudy (anonymous):

Well, its 9/8n = -54... Um, lets see you would multiply 9/8 to both sides first of all, so it would look like this 9/8*9/8n = -54 * 9/8 and then you would need to turn -54 into a fraction, so multiply 8 by -54 getting, 43/8 * 9/8, make sense ?

OpenStudy (anonymous):

yes mam thank you @I_Always_Smiling when u do 43/8*9/8 will u get your answer?

OpenStudy (anonymous):

Oh wait it's 432/8 * 9/8 sorry, typo!

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

@I_Always_Smiling so what would be your answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

You would have to multiply 432 by 9 and 8 by 8, then you would get your answer! Okay?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@I_Always_Smiling

OpenStudy (anonymous):

can u plz tell me the answer

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Oh yeah sure it would be 3888/ 16 simplified to 486/8.

OpenStudy (anonymous):

im only in 7th grade

OpenStudy (anonymous):

its hard as crap

OpenStudy (anonymous):

@I_Always_Smiling is 486 the answer?

OpenStudy (anonymous):

Which fractions are greater than -4/7 @i-am-always -smiling

OpenStudy (anonymous):

@just_one_last_goodbye